The probability that a student passes a class is

p(P) = 0.57.

The probability that a student studied for a class is

p(S) = 0.54.

The probability that a student passes a class given that he or she studied for the class is

p(P / S) = 0.74.

What is the probability that a student studied for the class, given that he or she passed the class

(p(S / P))?

Hint: Use Bayes' theorem. (Round your answer to two decimal places.)

p(S / P) =
Answer #1

P(P | S) = P(P and S) / P(S)

0.74 = P(P and S) / 0.54

P(P and S) = 0.74 * 0.54

= 0.3996

So, Using conditional Bayes' theorem,

P(S | P) = P( p and S ) / P(P)

= 0.3996 / 0.57

= 0.70